What is the value of QP for the reaction of nitrogen and oxygen to make dinitrogen tetroxide if the partial pressures of nitrogen is 1.2586 atm, the partial pressure of oxygen is 1.0884 atm and the partial pressure of dinitrogen tetroxide is 0.033054 atm?
Answer #1


The balanced equation is:

N2(g) + 2 O2(g) <-> N2O4(g)

Qp = p(N2O4) / p(N2)*p(O2)^2
= (0.033054) / (1.2586 * 1.0884^2)
= (0.033054) / (1.491)
= 2.217*10^-2

Answer: 2.217*10^-2